Of the 50 tables in a restaurant, 40% seat 2 people.
The remaining tables seat 4 people.
How many seats are there in the restaurant?


Answer by ikleyn(52776)

.

40%   of 50  tables  is/are  0.4*50 = 20 tables.


So, there are 20  tables with 2 seats and the rest  50-20 = 30 tables with 4 seats.


The total number of seats is 2*20 + 4*30 = 40 + 120 = 160.    ANSWER